OpenStudy (anonymous):

Evaluate the expression when x = 6. –2x – 4x –2 2 36 –36

pooja195 (pooja195):

-2(6)-4(6)

pooja195 (pooja195):

Whats -2 x 6=?

OpenStudy (anonymous):

-12

OpenStudy (peach3113):

What is 4 x 6=

pooja195 (pooja195):

^

OpenStudy (anonymous):

24

OpenStudy (peach3113):

-12-6=?

OpenStudy (peach3113):

Oh wait, sorry.

OpenStudy (peach3113):

-12-24=?

OpenStudy (anonymous):

so whats the answer

OpenStudy (peach3113):

24+12=?

OpenStudy (anonymous):

so whats the answer

pooja195 (pooja195):

u tell me

OpenStudy (anonymous):

i dont know what is it

pooja195 (pooja195):

the set up is wrong

pooja195 (pooja195):

-2(6)-4(6) -12+24=?

OpenStudy (anonymous):

no its not

OpenStudy (dayakar):

-2x-4x =-6x =-6*6 =-36

pooja195 (pooja195):

(−2)(6)−(4)(6) =−12−(4)(6) =−12−24 =?

OpenStudy (peach3113):

I forgot to say to make the answer negative after adding 24+12.

OpenStudy (peach3113):

My computer lagged and kicked me off.

OpenStudy (peach3113):

@juliannap If you need any more help, just tag me to your question.